# Customized SAM 2 for Referring Remote Sensing Image Segmentation

### v2

@@ Abstract
Referring Remote Sensing Image Segmentation (RRSIS) aims to segment target objects in remote sensing (RS) images based on textual descriptions. Although Segment Anything Model 2 (SAM 2) has shown remarkable performance in various segmentation tasks, its application to RRSIS presents several challenges, including understanding the text-described RS scenes and generating effective prompts from text descriptions. To address these issues, we propose RS2-SAM 2, a novel framework that adapts SAM 2 to RRSIS by aligning the adapted RS features and textual features, providing pseudo-mask-based dense prompts, and enforcing boundary constraints. Specifically, we first employ a union encoder to jointly encode the visual and textual inputs, generating aligned visual and text embeddings as well as multimodal class tokens. Then, we design a bidirectional hierarchical fusion module to adapt SAM 2 to RS scenes and align adapted visual features with the visually enhanced text embeddings, improving the model’s interpretation of text-described RS scenes. Additionally, a mask prompt generator is introduced to take the visual embeddings and class tokens as input and produce a pseudo-mask as the dense prompt of SAM 2. To further refine segmentation, we introduce a text-guided boundary loss to optimize segmentation boundaries by computing text-weighted gradient differences. Experimental results on several RRSIS benchmarks demonstrate that RS2-SAM 2 achieves state-of-the-art performance.
1 Introduction
@@
Referring Remote Sensing Image Segmentation (RRSIS) [19, 38] aims to segment specified targets from aerial images based on textual descriptions. This task extends the capabilities of traditional Referring Image Segmentation (RIS) [30, 35, 17] by addressing the unique challenges inherent to remote sensing images. These challenges include handling diverse spatial scales, interpreting complex scene contexts, and resolving ambiguous object boundaries, which are particularly prevalent in remote sensing scenes.
Recent advances in the Segment Anything Model (SAM) [10] and its variants [41, 32, 9] have demonstrated significant improvements in efficiency and accuracy for promptable segmentation tasks in natural images. These models exhibit powerful segmentation capabilities and robust interactive prompting mechanisms. Notably, SAM 2 [24] incorporates the hierarchical encoder Hiera [26], enhancing its ability to process images with diverse spatial scales. However, despite these advancements, applying SAM 2 to RRSIS remains challenging due to the unique complexities of remote sensing images.
First, while SAM 2 demonstrates strong performance on natural images, its effectiveness diminishes in remote sensing scenes due to the low distinguishability of targets and the limited contrast between foreground and background regions. Furthermore, SAM 2 struggles to effectively utilize spatial information derived from textual descriptions, as illustrated in fig. 1 (a). Although existing studies have attempted to address these issues, significant room for improvement remains. For example, SAM2-Adapter [1] modifies SAM 2’s encoding process to better adapt to complex scenes. However, this unimodal approach lacks hierarchical information interaction, resulting in an insufficient fine-grained understanding of textual information. Consequently, achieving effective alignment of vision-language information and improving generalization to complex remote sensing scenes remain critical challenges in adapting SAM 2 for RRSIS. Second, SAM 2’s inherent lack of textual prompt integration limits its ability to generate prompts that accurately align with given textual descriptions. Previous RRSIS methods, such as RMSIN [19], rely on independent encoding for visual and textual inputs, but such traditional encoder-decoder structures fail to effectively integrate textual information into SAM 2. EVF-SAM [43] addresses this by jointly encoding vision and language features and generating sparse prompts through a multilayer perceptron (MLP), achieving promising results on natural images. However, in remote sensing scenes, sparse prompts alone are insufficient for fine-grained local image understanding, particularly for recognizing subtle and less distinguishable objects. Thus, designing effective prompts based on text description to guide the decoding process remains a critical challenge for adapting SAM 2 to RRSIS. Additionally, objects in remote sensing images often blend with their surroundings, resulting in ambiguous boundaries. Enhancing the distinction between target objects and the background remains an urgent and unresolved issue.
@@
To address these challenges, this paper proposes RS2-SAM 2, a novel RRSIS framework adapted from SAM 2. As illustrated in fig. 1 (b), our approach focuses on three key aspects: (1) adapting SAM 2 image features to remote sensing scenes and aligning them with textual features during the image encoding process, (2) generating dense prompts for precise segmentation, and (3) enforcing boundary constraints throughout training. Specifically, to adapt SAM 2 to remote sensing scenes and align remote sensing visual features with textual features, we design a bidirectional hierarchical fusion module, which is embedded both within and after the frozen SAM 2 image encoder. Initially, we utilize a union encoder to jointly encode visual and textual inputs, producing semantically aligned visual and textual embeddings as well as multimodal class tokens. Subsequently, our proposed fusion module hierarchically aligns the visually enhanced textual embeddings with adapted remote sensing visual features in the SAM 2 encoder.
To equip SAM 2 with more precise prompts, we introduce a mask prompt generator. This module integrates jointly encoded visual embeddings and multimodal class tokens to generate a pseudo-mask representing remote sensing target objects. The pseudo-mask is then fed into the SAM 2 prompt encoder as a dense prompt, providing pixel-level positional information to enhance the model’s segmentation capability. Furthermore, we propose a text-guided boundary loss, which computes the gradient values of both the predicted and ground truth masks to better distinguish object boundaries. This loss is dynamically weighted by textual information, effectively constraining the predicted mask boundaries and improving segmentation accuracy.
@@
  • We introduce a text-guided boundary loss to refine segmentation accuracy by constraining boundary consistency.
@@
Figure 2: The overview of the proposed RS2-SAM 2 framework. It consists of four key components: the union encoder, the bidirectional hierarchical fusion module, the mask prompt generator, and SAM 2. The union encoder extracts multimodal representations from the input image and text. The bidirectional hierarchical fusion module enhances image features with textual embeddings. The mask prompt generator produces a prior mask as the dense prompt for SAM 2. Finally, SAM 2 generates precise segmentation masks, while the text-guided boundary loss constrains the boundary accuracy of the predicted masks.
2 Related work
@@
Referring Remote Sensing Image Segmentation. RRSIS emerges as an important research direction in remote sensing, allowing users to extract geospatial information through natural language queries. Compared to the remote sensing visual grounding (RSVG) [27, 40, 13] task, RRSIS focuses on fine-grained pixel-level analysis rather than region-level identification. However, research in this area remains in its early stages with limited exploration. Yuan et al. [38] pioneer this task by introducing the RefsegRS dataset and adapting the LAVT [35] framework for remote sensing. To address challenges like small and fragmented targets easily confused with complex backgrounds, they develop a multi-level feature fusion mechanism incorporating linguistic guidance, significantly enhancing small object detection. Subsequently, Liu et al. [19] build upon the RSVG dataset DIOR-RSVG [40] to establish the first large-scale RRSIS dataset, RRSIS-D. They further introduce the RMSIN architecture, which leverages rotational convolution operations to better handle spatial-scale variations and directional complexities in remote sensing images, providing a more robust modeling framework for RRSIS. Building upon RMSIN [19], Lei et al. propose FIANet [14], which focuses more on fine-grained vision-language interaction and adaptive understanding of objects at different scales.
@@
Segment Anything Model. SAM [10] is an interactive model designed to generate non-semantic segmentation masks based on a variety of input prompts. Leveraging large-scale training data, it demonstrates strong generalization across diverse segmentation tasks. SAM has also been widely utilized in multiple domains, including remote sensing image analysis [28], video object tracking [2, 34], and medical image segmentation [39, 3]. To improve its accuracy and efficiency, several optimized versions [41, 32, 44] have been developed. More recently, SAM 2 [24] has been introduced for video segmentation [25, 12], incorporating multi-scale feature encoding to enhance segmentation robustness. Despite these advances, SAM remains limited by its lack of linguistic understanding, preventing it from directly handling referring segmentation tasks that require textual guidance. To address this limitation, recent works [11, 33, 23, 42] have explored the integration of multimodal large language models (MLLMs) to enhance SAM’s ability to process language-based instructions. LISA [11], for example, fine-tunes LLaVA [16] to derive hidden embeddings, thereby producing multimodal features for enhanced segmentation. u-LLaVA [33] extends this paradigm by enabling simultaneous region- and pixel-level segmentation through multi-task learning. Due to its lightweight pre-fusion strategy, EVF-SAM [43] leverages joint visual-language encoding to enhance text-driven segmentation prompts, thereby achieving superior segmentation accuracy. However, while these models have demonstrated strong performance on natural images, their effectiveness in complex remote sensing scenes remains limited, highlighting the need for further refinement and adaptation.
3 Method
3.1 Overview
@@
The architecture of the proposed RS2-SAM 2 framework is depicted in fig. 2. RS2-SAM 2 is composed of four essential parts: the union encoder, the bidirectional hierarchical fusion module, the mask prompt generator, and the SAM 2 model. Given an input remote sensing image $\mathcal{I}$ and its associated textual description =$\mathcal{E}=\left\{e_{l}\right\}_{l=1}^{L}$, where $L$ represents the number of words, the union encoder processes both modalities to extract a multimodal [CLS] token, visual patch embeddings, and textual embeddings. The SAM 2 image encoder, which incorporates the bidirectional hierarchical fusion module, further refines the extracted image features by leveraging textual information. Next, the mask prompt generator utilizes the visual patch embeddings and the multimodal [CLS] token to produce a pseudo-mask estimate for the target object. This prior mask, along with the multimodal [CLS] token, acts as a guiding signal for SAM 2. Ultimately, the SAM 2 decoder synthesizes the extracted image features and generated prompts to produce high-precision segmentation masks, while the text-guided boundary loss constrains the accuracy of object contours.
3.2 Feature Extraction
For the RRSIS task, aligning the visual and linguistic modalities is vital. We adopt the union BEiT-3 [29] encoder to simultaneously process the visual and textual features, in line with the approaches in [43, 37]. Each input frame $I_{u}\in\mathbb{R}^{H_{u}\times W_{u}\times 3}$ is divided into non-overlapping patches $P_{v}\in\mathbb{R}^{N_{p}\times(p^{2}\times 3)}$, which are then projected into a $D$-dimensional space, yielding $P_{v}\in\mathbb{R}^{N_{p}\times D}$. Here, $N_{p}=\frac{H_{u}\times W_{u}}{p^{2}}$, and $D$ represents the embedding dimension. A visual class token $V_{cls}\in\mathbb{R}^{1\times D}$ is prepended, along with learnable positional embeddings $V_{pos}\in\mathbb{R}^{(N_{p}+1)\times D}$, to form the final visual representation $V_{0}\in\mathbb{R}^{(N_{p}+1)\times D}$. Concurrently, the textual input of length $L$ is tokenized using XLMRobertaTokenizer [4], yielding the tokenized sequence $T_{seq}$. A class token $T_{cls}\in\mathbb{R}^{1\times D}$ and end-of-sequence marker $T_{eos}\in\mathbb{R}^{1\times D}$ are appended, followed by positional embeddings $T_{pos}\in\mathbb{R}^{N_{t}\times D}$, where $N_{t}=L+2$. The final union multimodal representation $U_{0}$ is obtained by concatenating the visual and text embeddings, i.e., $U_{0}=\text{concat}(V_{0},T_{0})$. The above procedure is formally represented as:
$V_{0}=[V_{cls},P_{v}]+V_{pos},$ (1)
$T_{0}=[T_{cls},T_{seq},T_{eos}]+T_{pos},$
$U_{0}=[V_{0};T_{0}]\in\mathbb{R}^{(N_{p}+N_{t}+1)\times{D}}.$
After multimodal fusion through multiple attention blocks, the joint visual-text embedding is passed into modality-specific FFNs for both vision and text processing. Ultimately, we obtain the union visual-text embeddings $U\in{\mathbb{R}^{(N_{p}+N_{t}+1)\times{D}}}$, which are then decomposed into a multimodal [CLS] token $V_{cls}\in\mathbb{R}^{{1}\times{D}}$, visual patch embeddings $V\in\mathbb{R}^{{N_{p}}\times{D}}$, and text embeddings $T\in{\mathbb{R}^{{N_{t}}\times{D}}}$.
@@ 3.2 Feature Extraction
Subsequently, the SAM 2 image encoder, in conjunction with the proposed bidirectional hierarchical fusion module, encodes the input image $I_{s}\in\mathbb{R}^{H_{s}\times W_{s}\times 3}$. It extracts a set of hierarchical multi-scale features for mask production, while the final-layer image feature $F_{n}\in\mathbb{R}^{\frac{H_{s}}{16}\times\frac{W_{s}}{16}\times C}$ serves as the input for the subsequent decoding process. Here, $C$, $H_{s}$ and $W_{s}$ represent the channel number, height, and width of the SAM 2 image encoder, respectively.
3.3 Bidirectional Hierarchical Fusion Module
@@
Although SAM 2 demonstrates powerful segmentation capabilities for natural images, it struggles to achieve accurate segmentation for remote sensing images with more complex scenes. Inspired by MSA [31], a straightforward approach is to incorporate linear layers into the SAM 2 image encoder to enhance its adaptability to remote sensing images. However, for the RRSIS task, integrating textual information during SAM 2’s image encoding process could make the model more sensitive to referred objects. To address this, we designed a bidirectional hierarchical fusion module and embedded it into SAM 2 image encoder, enabling the visual feature of SAM 2 to better adapt to remote sensing scenes and the referring text, thereby achieving more precise segmentation.
@@
As illustrated in fig. 3, our bidirectional hierarchical fusion module begins with the preprocessed SAM 2 image feature $F_{i}$ of the current layer being projected to a lower dimensionality through a linear layer and an activation function, yielding $F_{i}^{{}^{\prime}}$. Simultaneously, the text feature of the current layer, $T_{i}$, is also projected via a linear layer to obtain $T_{i}^{{}^{\prime}}$, ensuring dimensional consistency with the image feature. Subsequently, the image and text features are utilized as query embeddings and key-value pairs, respectively, and undergo cross-attention interaction to capture modality-specific dependencies. The resulting features are then integrated with the pre-interaction representations through element-wise addition. Following this, linear layers are used to restore the dimensionality of both the image and text features, producing the updated representations $F_{i}^{{}^{\prime\prime}}$ and $T_{i}^{{}^{\prime\prime}}$.
@@
To preserve textual integrity during visual enhancement, the text feature is weighted and summed with its pre-interaction representation. Meanwhile, after skip-connection, the visual feature is summed with the original image feature $F_{in}$ of the current layer, then separately processed by the MLP branch and the linear branch, followed by weighted fusion. The above process can be expressed by the following equation:
@@
$F_{out}=$ $F^{{}^{\prime\prime\prime}}_{i}+\text{MLP}(\text{LN}(F^{{}^{\prime\prime\prime}}_{i}))+\alpha_{i}\text{Linear}(\sigma(\text{Linear}(F^{{}^{\prime\prime\prime}}_{i}))),$
@@
3.4 Mask Prompt Generator
@@
3.5 SAM 2 Prompt Encoder and Mask Decoder
The prompt encoder in RS2-SAM 2 accepts two types of inputs: the dense prompt $M_{p}$ derived from the mask prompt generator and the sparse prompt $V_{cls}$ produced by the union encoder. Following an approach similar to EVF-SAM [43], the [CLS] token $V_{cls}$ is passed through a token MLP and then combined with sparse embeddings initialized to zero. To ensure compatibility, the pixel-level prompt $M_{p}$ is adjusted to the spatial dimensions of SAM 2 features before being input as the dense spatial information into the mask decoder.
Within the SAM 2 mask decoder, both sparse object-level and dense pixel-level prompts are used. The sparse input is combined with object tokens to formulate queries for the decoder, while the dense pixel-level prompt, serving as the mask prior, is element-wise added to SAM 2 visual features to facilitate direct decoding from high-resolution feature representations.
3.6 Training Loss
RS2-SAM 2 utilizes a comprehensive loss function akin to that proposed in [19] to regulate the predicted mask, formulated as follows:
@@ 3.6 Training Loss
where $\mathcal{L}_{ce}$ corresponds to the cross-entropy loss, $\mathcal{L}_{dice}$ denotes the DICE loss [22], and $\mathcal{L}_{tbl}$ signifies the text-guided boundary loss, a loss function we designed that uses text-weighted constraints to predict mask boundaries.
@@
$\nabla_{pre}=Absd_{h}(M_{pre})+Absd_{v}(M_{pre}),$ (4)
$\nabla_{gt}=Absd_{h}(M_{gt})+Absd_{v}(M_{gt}),$
$\mathcal{L}_{tbl}=\frac{1}{N}\sum_{i=1}^{N}(\text{Linear}(T_{s})({\nabla_{pre}}_{i}-{\nabla_{gt}}_{i}))^{2},$
@@
4 Experiments
4.1 Datasets and Metrics
@@
Datasets. The experiments are performed on two key RRSIS datasets: RefSegRS [38] and RRSIS-D [19]. The RefSegRS dataset is the first remote sensing referring segmentation dataset, consisting of 2172 images for training, 413 images for validation, and 1817 images for testing. Each image in the dataset has a resolution of 512×512. RRSIS-D is a widely recognized and large-scale dataset in the field of RRSIS, containing 12181 samples in the training set, 1740 samples in the validation set and 3481 samples in the test set. The resolution of each image is 800×800.
@@
Evaluation Metrics. In accordance with the evaluation protocol established in [19], we assess our model using several metrics, including precision at various IoU thresholds (Pr@0.5 to Pr@0.9), mean Intersection-over-Union (mIoU), and overall Intersection-over-Union (oIoU). These metrics are computed on the validation and test sets of both RefSegRS and RRSIS-D datasets.
4.2 Implementation Details

### v3

# RS2-SAM2: Customized SAM2 for Referring Remote Sensing Image Segmentation

@@ Abstract
Referring Remote Sensing Image Segmentation (RRSIS) aims to segment target objects in remote sensing (RS) images based on textual descriptions. Although Segment Anything Model 2 (SAM2) has shown remarkable performance in various segmentation tasks, its application to RRSIS presents several challenges, including understanding the text-described RS scenes and generating effective prompts from text descriptions. To address these issues, we propose RS2-SAM2, a novel framework that adapts SAM2 to RRSIS by aligning the adapted RS features and textual features, providing pseudo-mask-based dense prompts, and enforcing boundary constraints. Specifically, we employ a union encoder to jointly encode the visual and textual inputs, generating aligned visual and text embeddings as well as multimodal class tokens. A bidirectional hierarchical fusion module is introduced to adapt SAM2 to RS scenes and align adapted visual features with the visually enhanced text embeddings, improving the model’s interpretation of text-described RS scenes. To provide precise target cues for SAM2, we design a mask prompt generator, which takes the visual embeddings and class tokens as input and produces a pseudo-mask as the dense prompt of SAM2. Experimental results on several RRSIS benchmarks demonstrate that RS2-SAM2 achieves state-of-the-art performance.
Introduction
@@
Referring Remote Sensing Image Segmentation (RRSIS) (Liu et al. 2024b; Yuan et al. 2024) aims to segment specified targets from aerial images based on textual descriptions. This task extends the capabilities of traditional Referring Image Segmentation (RIS) (Wang et al. 2022; Yang et al. 2022; Liu et al. 2023a) by addressing the unique challenges inherent to remote sensing images. These challenges include handling diverse spatial scales, interpreting complex scene contexts, and resolving ambiguous object boundaries, which are particularly prevalent in remote sensing scenes.
@@
Recent advances in the Segment Anything Model (SAM) (Kirillov et al. 2023) and its variants (Zhang et al. 2023; Xiong et al. 2024; Ke et al. 2024) have demonstrated significant improvements in efficiency and accuracy for promptable segmentation tasks in natural images. These models exhibit powerful segmentation capabilities and robust interactive prompting mechanisms. Notably, SAM2 (Ravi et al. 2024) incorporates the hierarchical encoder Hiera (Ryali et al. 2023), enhancing its ability to process images with diverse spatial scales. However, despite these advancements, applying SAM2 to RRSIS remains challenging due to the unique complexities of remote sensing images.
First, while SAM2 demonstrates strong performance on natural images, its effectiveness diminishes in remote sensing scenes due to the low distinguishability of targets and the limited contrast between foreground and background regions. Furthermore, SAM2 struggles to effectively utilize spatial information derived from textual descriptions, as illustrated in Fig. 1 (a). Although existing studies have attempted to address these issues, significant room for improvement remains. For example, SAM2-Adapter (Chen et al. 2024) modifies SAM2’s encoding process to better adapt to complex scenes. However, this unimodal approach lacks hierarchical information interaction, resulting in an insufficient fine-grained understanding of textual information. Consequently, achieving effective alignment of vision-language information and improving generalization to complex remote sensing scenes remain critical challenges in adapting SAM2 for RRSIS. Second, SAM2’s inherent lack of textual prompt integration limits its ability to generate prompts that accurately align with given textual descriptions. Previous RRSIS methods, such as RMSIN (Liu et al. 2024b), rely on independent encoding for visual and textual inputs, but such traditional encoder-decoder structures fail to effectively integrate textual information into SAM2. EVF-SAM (Zhang et al. 2024) addresses this by jointly encoding vision and language features and generating sparse prompts through a multilayer perceptron (MLP), achieving promising results on natural images. However, in remote sensing scenes, sparse prompts alone are insufficient for fine-grained local image understanding, particularly for recognizing subtle and less distinguishable objects. Thus, designing effective prompts based on text description to guide the decoding process remains a critical challenge for adapting SAM2 to RRSIS.
To address these challenges, this paper proposes RS2-SAM2, a novel RRSIS framework adapted from SAM2. As illustrated in Fig. 1 (b), our approach focuses on two key aspects: (1) adapting SAM2 image features to remote sensing scenes and aligning them with textual features during the image encoding process, and (2) generating dense prompts for precise segmentation. Specifically, to adapt SAM2 to remote sensing scenes and align remote sensing visual features with textual features, we design a bidirectional hierarchical fusion module, which is embedded both within and after the frozen SAM2 image encoder. Initially, we utilize a union encoder to jointly encode visual and textual inputs, producing semantically aligned visual and textual embeddings as well as multimodal class tokens. Subsequently, our proposed fusion module hierarchically aligns the visually enhanced textual embeddings with adapted remote sensing visual features in the SAM2 encoder.
To equip SAM2 with more precise prompts, we introduce a mask prompt generator. This module integrates jointly encoded visual embeddings and multimodal class tokens to generate a pseudo-mask representing remote sensing target objects. The pseudo-mask is then fed into the SAM2 prompt encoder as a dense prompt, providing pixel-level positional information to enhance the model’s segmentation capability.
@@
Figure 2: The overview of the proposed RS2-SAM2 framework. It consists of four key components: the union encoder, the bidirectional hierarchical fusion module, the mask prompt generator, and SAM2. The union encoder extracts multimodal representations from the input image and text. The bidirectional hierarchical fusion module enhances image features with textual embeddings. The mask prompt generator produces a prior mask as the dense prompt for SAM2. Finally, SAM2 generates precise masks, while the text-guided boundary loss constrains their boundary accuracy.
Related work
@@
Referring Remote Sensing Image Segmentation. RRSIS emerges as an important research direction in remote sensing, allowing users to extract geospatial information through natural language queries. Compared to the remote sensing visual grounding (RSVG) (Sun et al. 2022; Zhan, Xiong, and Yuan 2023; Lan et al. 2024) task, RRSIS focuses on fine-grained pixel-level analysis rather than region-level identification. However, research in this area remains in its early stages with limited exploration. Yuan et al. (Yuan et al. 2024) pioneer this task by introducing the RefsegRS dataset and adapting the LAVT (Yang et al. 2022) framework for remote sensing. To address challenges like small and fragmented targets easily confused with complex backgrounds, they develop a multi-level feature fusion mechanism incorporating linguistic guidance, significantly enhancing small object detection. Subsequently, Liu et al. (Liu et al. 2024b) build upon the RSVG dataset DIOR-RSVG (Zhan, Xiong, and Yuan 2023) to establish the first large-scale RRSIS dataset, RRSIS-D. They further introduce the RMSIN architecture, which leverages rotational convolution operations to better handle spatial-scale variations and directional complexities in remote sensing images, providing a more robust modeling framework for RRSIS. Building upon RMSIN (Liu et al. 2024b), Lei et al. propose FIANet (Lei et al. 2024), which focuses more on fine-grained vision-language interaction and adaptive understanding of objects at different scales.
@@
Segment Anything Model. SAM (Kirillov et al. 2023) is an interactive model designed to generate non-semantic segmentation masks based on a variety of input prompts. Leveraging large-scale training data, it demonstrates strong generalization across diverse segmentation tasks. SAM has also been widely utilized in multiple domains, including remote sensing image analysis (Wang et al. 2024), video object tracking (Cheng et al. 2023; Yan et al. 2024), and medical image segmentation (Yue et al. 2024; Cheng et al. 2024). To improve its accuracy and efficiency, several optimized versions (Zhang et al. 2023; Xiong et al. 2024; Zhong et al. 2024) have been developed. More recently, SAM2 (Ravi et al. 2024) has been introduced for video segmentation (Rong et al. 2025; Lan et al. 2023), incorporating multi-scale feature encoding to enhance segmentation robustness. Despite these advances, SAM remains limited by its lack of linguistic understanding, preventing it from directly handling referring segmentation tasks that require textual guidance. To address this limitation, recent works (Lai et al. 2024; Xu et al. 2023; Rasheed et al. 2024) have explored the integration of multimodal large language models (MLLMs) to enhance SAM’s ability to process language-based instructions. LISA (Lai et al. 2024), for example, fine-tunes LLaVA (Liu et al. 2024a) to derive hidden embeddings, thereby producing multimodal features for enhanced segmentation. u-LLaVA (Xu et al. 2023) extends this paradigm by enabling simultaneous region- and pixel-level segmentation through multi-task learning. Due to its lightweight pre-fusion strategy, EVF-SAM (Zhang et al. 2024) leverages joint visual-language encoding to enhance text-driven segmentation prompts, thereby achieving superior segmentation accuracy. However, while these models have demonstrated strong performance on natural images, their effectiveness in complex remote sensing scenes remains limited, highlighting the need for further refinement and adaptation.
Method
Overview
@@
The architecture of the proposed RS2-SAM2 framework is depicted in Fig. 2. RS2-SAM2 comprises four essential parts: the union encoder, the bidirectional hierarchical fusion module, the mask prompt generator, and the SAM2 model. Given an input remote sensing image $\mathcal{I}$ and its associated textual description =$\mathcal{E}=\left\{e_{l}\right\}_{l=1}^{L}$, where $L$ is the number of words, the union encoder processes both modalities to extract a multimodal [CLS] token, visual patch embeddings, and textual embeddings. The SAM2 image encoder, incorporating the bidirectional hierarchical fusion module, further refines the extracted image features by leveraging textual information. Next, the mask prompt generator utilizes the visual patch embeddings and the multimodal [CLS] token to produce a pseudo-mask estimate for the target object. This prior mask, along with the multimodal [CLS] token, serves as a guiding signal for SAM2. Ultimately, the SAM2 decoder synthesizes the extracted image features and generated prompts to produce high-precision segmentation masks.
Feature Extraction
To align visual and textual modalities in RRSIS, this work uses the union BEiT-3 encoder (Wang et al. 2023), following (Zhang et al. 2024; Yu et al. 2024). Each input image $I_{u}\in\mathbb{R}^{H_{u}\times W_{u}\times 3}$ is divided into non-overlapping patches $P_{v}\in\mathbb{R}^{N_{p}\times(p^{2}\times 3)}$ and then projected to $P_{v}\in\mathbb{R}^{N_{p}\times D}$, where $N_{p}=\frac{H_{u}\times W_{u}}{p^{2}}$ and $D$ is the embedding dimension. A visual class token $V_{\text{cls}}\in\mathbb{R}^{1\times D}$ and positional embedding $V_{\text{pos}}\in\mathbb{R}^{(N_{p}+1)\times D}$ are prepended to obtain:
$V_{0}=[V_{\text{cls}},P_{v}]+V_{\text{pos}}$
Text of length $L$ is tokenized using XLM-Roberta (Conneau 2019) as $T_{\text{seq}}$. A class token and end-of-sequence marker are added, followed by positional embedding: $T_{0}=[T_{\text{cls}},T_{\text{seq}},T_{\text{eos}}]+T_{\text{pos}}$, where $T_{0}\in\mathbb{R}^{N_{t}\times D}$ and $N_{t}=L+2$.
The concatenated multimodal representation $U_{0}$ is constructed as:
$U_{0}=[V_{0};T_{0}]\in\mathbb{R}^{(N_{p}+N_{t}+1)\times D}$
After multimodal fusion and FFNs, the final representation $U$ is obtained, which is decomposed into $V_{\text{cls}}\in\mathbb{R}^{1\times D}$, $V\in\mathbb{R}^{N_{p}\times D}$, and $T\in\mathbb{R}^{N_{t}\times D}$.
@@
The SAM2 image encoder with the proposed bidirectional hierarchical fusion module encodes the input image $I_{s}\in\mathbb{R}^{H_{s}\times W_{s}\times 3}$. It extracts multi-scale features, and the final-layer output $F_{n}\in\mathbb{R}^{\frac{H_{s}}{16}\times\frac{W_{s}}{16}\times C}$ is used for decoding, where $H_{s}$, $W_{s}$, and $C$ denote the height, width, and channel number of the SAM2 image encoder, respectively.
Bidirectional Hierarchical Fusion Module
@@
Although SAM2 demonstrates powerful segmentation capabilities for natural images, it struggles to achieve accurate segmentation for remote sensing images with more complex scenes. Inspired by MSA (Wu et al. 2023), a straightforward approach is to incorporate linear layers into the SAM2 image encoder to enhance its adaptability to remote sensing images. However, for the RRSIS task, integrating textual information during SAM2’s image encoding process could make the model more sensitive to referred objects. To address this, we designed a bidirectional hierarchical fusion module and embedded it into SAM2 image encoder, enabling the visual feature of SAM2 to better adapt to remote sensing scenes and the referring text, thereby achieving more precise segmentation.
@@
As illustrated in Fig. 3, our bidirectional hierarchical fusion module begins with the preprocessed SAM2 image feature $F_{i}$ of the current layer being projected to a lower dimensionality through a linear layer and an activation function. Simultaneously, the text feature $T_{i}$ of the current layer is also projected to match the dimensionality of the image feature using a linear layer. Subsequently, the image and text features are utilized as query embeddings and key-value pairs, respectively, and undergo cross-attention interaction to capture modality-specific dependencies. The resulting features are then integrated with the pre-interaction representations through element-wise addition. Following this, linear layers are used to restore the dimensionality of both the image and text features. The above process can be expressed by the following equation:
$F_{i}^{{}^{\prime}}=\sigma(\text{Linear}(F_{i})),T_{i}^{{}^{\prime}}=\text{Linear}(T_{i}),$ (1)
$F_{i}^{{}^{\prime\prime}}=\text{MHCA}(F_{i}^{{}^{\prime}},T_{i}^{{}^{\prime}})+F^{{}^{\prime}}_{i},$
$T_{i}^{{}^{\prime\prime}}=\text{MHCA}(T_{i}^{{}^{\prime}},F_{i}^{{}^{\prime}})+T^{{}^{\prime}}_{i},$
where $\sigma$ denotes the activation function GeLU and MHCA represents the multi-head cross-attention layer.
@@
To preserve textual integrity during visual enhancement, the text feature is weighted and summed with its pre-interaction representation. Meanwhile, after skip-connection, the visual feature is summed with the original image feature $F_{in}$ of the current layer, then separately processed by the MLP branch and the linear branch, followed by weighted fusion. The entire process is depicted as follows:
@@
$F_{out}=$ $F^{{}^{\prime\prime\prime}}_{i}+\text{MLP}(\text{LN}(F^{{}^{\prime\prime\prime}}_{i}))+$
$\alpha_{i}\text{Linear}(\sigma(\text{Linear}(F^{{}^{\prime\prime\prime}}_{i}))),$
@@
Mask Prompt Generator
@@
SAM2 Prompt Encoder and Mask Decoder
The prompt encoder in RS2-SAM2 receives two inputs: a dense prompt $M_{p}$ from the mask prompt generator and a sparse prompt $V_{\text{cls}}$ from the union encoder. Following EVF-SAM (Zhang et al. 2024), $V_{\text{cls}}$ is processed via a token MLP and combined with zero-initialized sparse embeddings. To ensure compatibility, the pixel-level prompt $M_{p}$ is adjusted to the spatial dimensions of SAM2 features before being input as the dense spatial information into the mask decoder.
The SAM2 mask decoder leverages both sparse object-level and dense pixel-level prompts: the sparse forms queries with object tokens, while the latter, as a mask prior, is added to visual features to enable direct high-resolution decoding.
Training Loss
RS2-SAM2 utilizes a comprehensive loss function akin to that proposed in (Liu et al. 2024b) to regulate the predicted mask, formulated as follows:
@@
where $\mathcal{L}_{ce}$ corresponds to the cross-entropy loss, $\mathcal{L}_{dice}$ denotes the DICE loss (Milletari, Navab, and Ahmadi 2016), and $\mathcal{L}_{tbl}$ signifies the text-guided boundary loss, a loss function we designed that uses text-weighted constraints to predict mask boundaries.
@@
$\nabla_{pre}=Absd_{h}(M_{pre})+Absd_{v}(M_{pre})$ (4)
$\nabla_{gt}=Absd_{h}(M_{gt})+Absd_{v}(M_{gt})$
$\mathcal{L}_{tbl}=\frac{1}{N}\sum_{i=1}^{N}(\text{Linear}(T_{s})({\nabla_{pre}}_{i}-{\nabla_{gt}}_{i}))^{2}$
@@
Experiments
Datasets and Metrics
@@
Datasets. The experiments are conducted on two key RRSIS datasets: RefSegRS (Yuan et al. 2024) and RRSIS-D (Liu et al. 2024b). RefSegRS is the first remote sensing referring segmentation dataset, containing 2172 images for training, 413 for validation, and 1817 for testing, each with a resolution of 512×512. RRSIS-D is a widely used large-scale dataset in the RRSIS field, including 12181 training samples, 1740 validation samples, and 3481 test samples, with each image sized at 800×800.
@@
Evaluation Metrics. In accordance with the evaluation protocol established in (Liu et al. 2024b), we assess our model using several metrics, including precision at various IoU thresholds (Pr@0.5 to Pr@0.9), mean Intersection-over-Union (mIoU), and overall Intersection-over-Union (oIoU). These metrics are computed on the validation and test sets of both RefSegRS and RRSIS-D datasets.
Implementation Details